# DCT-Based Decorrelated Attention for Vision Transformers

### v2

@@ 3.4 Frequency Domain Motivation of the DCT-based Initialization
for $n=0,1,\ldots,C-1$. The basis member $\cos\frac{(2n+1)k\pi}{2nC}$ is the $k$-th Chebyshev polynomial $T_{k}[\xi]$ evaluated at the $n$-th zero of $T_{n}[\xi]$ [1]. Figure 2 draws the basis vectors in Eq. (7) for an $8\times 8$ DCT matrix and the magnitude of their discrete Fourier transform (${}\mathcal{F}\{\cdot\}$):
@@
(a) $D_{0}$.
(b) $D_{1}$.
(c) $D_{2}$.
(d) $D_{3}$.
(e) $D_{4}$.
(f) $D_{5}$.
(g) $D_{6}$.
(h) $D_{7}$.
(i) Frequency response magnitudes. Frequencies are normalized into $[0,\pi]$.
Figure 2: Basis vectors of an $8\times 8$ DCT matrix and their frequency response magnitudes. $D_{i}$ refers to the $i$-th column. The basis vectors provide a good approximation to the eigenvectors of the Teoplitz matrix with $\rho=0.9$ (KLT). Each frequency response is shown in a solid curve in different colors, and their squared sum is denoted in the black dashed line, showing the entire spectrum is covered.
3.5 Motivation of the DCT-based Compression
The mechanism behind the DCT-based compression mimics the Karhunen-Loeve transform (KLT), which is also known as principal component analysis (PCA), a mathematical technique used for dimensionality reduction in signal processing [24, 61]. Neighboring image patches are highly correlated with each other. During training, the attention matrices are inter-correlated because they are obtained using similar gradient vectors. Therefore, they can be decorrelated and compressed using DCT. When the input data is highly correlated, the covariance matrix can be approximated by a Teoplitz matrix $\phi$ of $[1,\rho,\rho^{2},\cdots,\rho^{C-1}]$ with $\rho<1$ close to 1 [2]. This is the covariance matrix of an autoregressive model random process AR(1) with the correlation coefficient $\rho$ [22, 46]. Figure 2 shows that DCT basis vectors approximate the KLT of a first-order Markov process with high correlation among the neighboring samples [13, 38]. We observe that $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ have parameters with high correlations among themselves. We can directly use DCT in data compression since its matrix approximates the corresponding KLT matrix [1],
@@ 3.5 Motivation of the DCT-based Compression
4 Experimental Results
We conduct experiments on the CIFAR-10 [25] and ImageNet-1K [12] classification tasks and COCO 2017 detection and segmentation tasks [28]. Swin Transformers [29] are employed as the backbone network. Implementation details are presented in the technical appendices.
4.1 CIFAR-10 Classification
Training vision transformers from scratch on the CIFAR-10 dataset is notoriously difficult. We demonstrate the benefits of using our DCT initialization and DCT compression for the Swin Transformer Tiny (Swin-T) on from-scratch CIFAR-10 training to show the effectiveness of our methods on the small-scale dataset. Table 2 presents the performance comparison on the CIFAR-10 test dataset, where the number of parameters only counts trainable parameters.
Table 2: CIFAR-10 test dataset evaluation. In column “DCT", Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Models with higher or lower accuracy are highlighted in pink or cyan, respectively.
@@ 4.1 CIFAR-10 Classification
Next, we evaluate the DCT-compressed attention with different truncating rates in Swin-T on the CIFAR-10 dataset. Table 2 presents when we only keep 25% of DCT coefficients, the number of trainable parameters is reduced from 27.53M to 21.45M (22.1% off), while the accuracy only drops 0.05%. If we keep more DCT coefficients, the accuracy even improves. If we keep 75% of DCT coefficients, the accuracy increases from 90.51% to 90.90%. This is because although the number of trainable parameters is reduced, transforming the input data into the frequency domain using DCT can boost the extraction of important features. As a result, fewer trainable parameters are required.
4.2 ImageNet-1K Classification
Table 3 presents the performance comparison on the ImageNet-1K dataset, where the number of parameters only counts trainable parameters. First, we investigate the DCT-initialized attention in the Swin-T and the Swin-S on the ImageNet-1K dataset. The accuracy on the ImageNet-1K validation dataset is always improved nevertheless which weight matrix is initialized as the DCT matrix. For example, when we initialize the weights to compute keys ($\mathbf{K}$) as the DCT weights, the accuracy of Swin-T from 81.474% to 81.728% and Swin-S are improved from 83.196% to 83.378%, respectively.
Table 3: ImageNet-1K validation dataset evaluation. After DCT, Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Our models with higher or lower top-1 accuracy are highlighted in pink or cyan, respectively.
@@ 4.2 ImageNet-1K Classification
Then, we evaluate the DCT-compressed attention with different truncating rates and compare the results with DCFormer [26], as it is also a DCT-based method to reduce the computational cost. Table 3 presents that when we keep 75% of DCT coefficients, we can obtain a comparable accuracy result on Swin-T and a slightly better accuracy result on Swin-S. Contrarily, the trainable parameters of the Swin-T and the Swin-S are reduced from 28.29M to 25.45M (13.1% off) and from 49.61M to 44.45M (10.4% off), respectively. Flops are reduced from 4.49G to 4.18G (6.9% off) and 8.13G to 8.74G (7.0% off), respectively. If fewer DCT coefficients are kept, the swim transformers can get further savings on parameters and computational overhead. Meanwhile, the accuracy will drop correspondingly. As a comparison, although DCFormer-SW-T/S [26] with $\tau=2$ reduce the FLOPs significantly ($\tau=2$ means 4×4$\times$4 × frequency compression), the performance drops seriously because the DCFormer takes both low-pass and high-pass DCT components equivalent important.
4.3 Ablation Study
Multiple DCT Initialization Table 4 presents the ablation study about initializing multiple matrices of $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ as the DCT matrix. Although such an initialization still improves the accuracy for Swin-T and Swin-S, the improvement is smaller than with the single DCT initialization. For example, in the Swin-T experiments, if $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$ are both initialized as the DCT matrix, the top-1 accuracy is 81.540%, lower than 81.568% from initializing $\mathbf{W}_{Q}$ only and 81.728% from initializing $\mathbf{W}_{K}$ only. This deterioration is caused by the weight symmetry problem [20] from initializing the multiple weights as the same values. In other words, the attention mechanism is designed to attend to relevant information in the input sequence. If $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the same, the model may not effectively capture the relationships between different positions in the input sequence. Moreover, one of the advantages of self-attention is its ability to model complex dependencies between different positions in the sequence. If $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are the same, the capacity of the model to capture diverse patterns will be limited, and the performance of the model will drop correspondingly.
Table 4: ImageNet-1K ablation study on multiple weights initialization as the DCT matrix.
@@ 4.3 Ablation Study
Non-Trainable DCT Weights Initialization In this case, computing queries, keys, and values is equivalent to shifting the DCT representation of the input patterns. Table 5 shows we obtain comparable accuracy results as the baseline but use fewer trainable parameters on the CIFAR-10 dataset. It also shows on the ImageNet-1K dataset if we take non-trainable DCT weights to compute $\mathbf{Q}$ or $\mathbf{K}$, the top-1 accuracy drops insignificantly (less than 0.3%). It is worthwhile noting that in this work, the DCT is implemented by matrix multiplication. Hence, the computational overhead of using non-trainable DCT weights is the same as the vanilla attention. Still, the DCT with butterfly operation can further reduce this cost as we only need to apply DCT in each attention layer.
Table 5: Ablation study on non-trainable DCT weights.
@@
4.4 COCO 2017 Object Detection and Instance Segmentation
We further evaluate the proposed methods on COCO 2017 object detection and instance segmentation tasks [28]. We use the Swin Transformers with ImageNet-1K pre-trained weights as the backbone and the Cascade Mask R-CNN [6, 7] as the pipeline. Since initializing $\mathbf{W}_{K}$ as the DCT matrix obtains the best accuracy results in the ImageNet-1K experiments, we use Swin-T-DCT-K and Swin-S-DCT-K to evaluate the DCT-based initialization. As Table 6 presents, both Swin-T-DCT-K and Swin-S-DCT-K obtain slightly better average precision (AP) results than the vanilla models in the two tasks. Swin-T-DCT-0.75 obtains comparable AP results as Swin-T with about 3M fewer parameters, and Swin-S-DCT-0.75 gets better AP results than Swin-S with about 5M fewer parameters.
Table 6: COCO results. Our methods with higher or lower $\text{AP}^{\text{Mask}}$ are highlighted in pink or cyan.
@@ 4.4 COCO 2017 Object Detection and Instance Segmentation
5 Limitation
DCT in this work is implemented by matrix multiplication because the fast DCT is not officially supported in PyTorch. In the future, we will implement fast DCT to further simplify the computation.
6 Conclusion
In this work, we introduced two new methods using Discrete Cosine Transform (DCT) to decorrelate the input in the attention for vision transformers. First, we presented a novel DCT-based initialization approach to address the challenges associated with training attention weights. We demonstrated a significant improvement over traditional initialization, providing a robust foundation for the attention mechanism. Second, we formulated a DCT-based compression method that can effectively reduce the computational overhead by truncating high-frequency components. It resulted in smaller weight matrices with compromising considerable accuracy. We validated the efficacy of these breakthroughs, showcasing their potential to enhance the performance and efficiency of Transformer models in the classification tasks. These contributions underscore the importance of thoughtful initialization strategies and compression techniques in advancing the capabilities of the Swin Transformer architectures.
Appendix A Technical Appendices
A.1 Implementation Details
Models are trained on a server with 8 NVIDIA A100 80GB GPUs using Python with PyTorch. The code of this work will be released upon the paper’s acceptance.
A.1.1 Classification Tasks Implementation Details
Training models for the image classification tasks from scratch mostly follow the settings in [29]. We employ AdamW optimizer [30] for 300 epochs using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We set the batch size as 1024, the initial learning rate as 0.001, and the weight decay as 0.05. Data augmentation is performed as the following: The CIFAR-10 augmentation follows [19]. First, the training images are padded with 4 pixels. Then, they are randomly cropped to get 32 by 32. Finally, the images are randomly flipped horizontally. The images are normalized with the means of [0.4914, 0.4822, 0.4465] and the standard variations of [0.2023, 0.1994, 02010]. The ImageNet-1K augmentation follows PyTorch Torchvision’s GitHub repository [37] [^1] https://github.com/pytorch/vision/tree/release/2.0/references/classification. As a start, images are resized to make the smaller edge of the image 232 for Swin-T and 246 for Swin-S. Next, they are cropped to 224 by 224 images and randomly flipped horizontally. Techniques such as Autoaugment [11], mixup [58], cutmix [56], and random erasing [63] are applied. Images are normalized with the means of [0.485, 0.456, 0.406] and the standard variations of [0.229, 0.224, 0.225]. During the training, the best models are stored based on the top-1 accuracy on the CIFAR-10 test dataset and the ImageNet-1K validation dataset.
Flops computation for the Swin Transformers and our revisions uses the function “flops” in the official implementation of [29] [^2] https://github.com/microsoft/Swin-Transformer/blob/main/models/swin_transformer.py. To follow their work, the flops from the Softmax function are omitted.
Flops computation for the ViT-B-32 and our revisions uses the function “ptflops” in the official implementation of [44], with manually adding the flops from DCT.
A.1.2 COCO Detection and Segmentation Tasks Implementation Details
Training models for the COCO tasks follow the settings in [29]. Models start from ImageNet-1K pre-trained weights. We use AdamW optimizer [30] for 36 epochs with an initial learning rate of 0.0001, weight decay of 0.05, batch size of 16, 3$\times$ schedule, and multi-scale training [8, 45].

### v3

@@ III-D Frequency Domain Motivation of the DCT-based Initialization
for $n=0,1,\ldots,C-1$. The basis member $\cos\frac{(2n+1)k\pi}{2nC}$ is the $k$-th Chebyshev polynomial $T_{k}[\xi]$ evaluated at the $n$-th zero of $T_{n}[\xi]$ [41]. Fig. 2 draws the basis vectors in Eq. (7) for an $8\times 8$ DCT matrix and Fig. 3 presents the magnitude of their discrete Fourier transform (${}\mathcal{F}\{\cdot\}$):
@@
(a) $D_{0}$
(b) $D_{1}$
(c) $D_{2}$
(d) $D_{3}$
(e) $D_{4}$
(f) $D_{5}$
(g) $D_{6}$
(h) $D_{7}$
Figure 2: Basis vectors of an $8\times 8$ DCT matrix where $D_{i}$ refers to the $i$-th column. The DCT basis vectors provide a good approximation to the eigenvectors of the Teoplitz matrix with $\rho=0.9$ (KLT).
Figure 3: Frequency response of the DCT basis vectors. Each frequency response is shown in a solid curve in different colors. The frequency is normalized to $[0,\pi]$. The black dashed line denotes the root of the squared sum of these frequency responses, showing the entire spectrum is covered.
III-E Motivation of the DCT-based Compression
The mechanism behind the DCT-based compression mimics the Karhunen-Loeve transform (KLT), which is also known as principal component analysis (PCA), a mathematical technique used for dimensionality reduction in signal processing [43, 44]. Neighboring image patches are highly correlated with each other. During training, the attention matrices are inter-correlated because they are obtained using similar gradient vectors. Therefore, they can be decorrelated and compressed using DCT. When the input data is highly correlated, the covariance matrix can be approximated by a Teoplitz matrix $\phi$ of $[1,\rho,\rho^{2},\cdots,\rho^{C-1}]$:
$\phi=\begin{bmatrix}1&\rho&\rho^{2}&\cdots&\rho^{C-1}\cr\rho&1&\rho&\cdots&\rho^{C-2}\cr\rho^{2}&\rho&1&\cdots&\rho^{C-3}\cr\vdots&\vdots&\vdots&\ddots&\vdots\cr\rho^{C-1}&\rho^{C-2}&\rho^{C-3}&\cdots&1\cr\end{bmatrix},$ (9)
with $\rho<1$ close to 1 [45]. This is the covariance matrix of an autoregressive model random process AR(1) with the correlation coefficient $\rho$ [46, 47]. Fig. 2 shows that DCT basis vectors approximate the KLT of a first-order Markov process with high correlation among the neighboring samples [48, 49]. We observe that $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ have parameters with high correlations among themselves. We can directly use DCT in data compression since its matrix approximates the corresponding KLT matrix [41],
@@ III-E Motivation of the DCT-based Compression
IV Experimental Results
We conduct experiments on the CIFAR-10 [53] and ImageNet-1K [54] classification tasks and COCO 2017 detection and segmentation tasks [55]. Swin Transformers [17] are employed as the backbone network. Models are trained on a server with 8 NVIDIA A100 80GB GPUs using Python with PyTorch framework [56].
IV-A Implementation Details
IV-A1 Image Classification Tasks Implementation Details
Training models for the image classification tasks from scratch mostly follow the settings in [17]. We employ AdamW optimizer [57] for 300 epochs using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We set the batch size as 1024, the initial learning rate as 0.001, and the weight decay as 0.05. Data augmentation is performed as the following: The CIFAR-10 augmentation follows [58]. First, the training images are padded with 4 pixels. Then, they are randomly cropped to get 32 by 32. Finally, the images are randomly flipped horizontally. The images are normalized with the means of [0.4914, 0.4822, 0.4465] and the standard variations of [0.2023, 0.1994, 02010]. The ImageNet-1K augmentation follows PyTorch Torchvision’s GitHub repository [56] [^1] https://github.com/pytorch/vision/tree/release/2.0/references/classification. As a start, images are resized to make the smaller edge of the image 232 for Swin-T and 246 for Swin-S. Next, they are cropped to 224 by 224 images and randomly flipped horizontally. Techniques such as Autoaugment [59], mixup [60], cutmix [61], and random erasing [62] are applied. Images are normalized with the means of [0.485, 0.456, 0.406] and the standard deviations of [0.229, 0.224, 0.225]. During the training, the best models are stored based on the top-1 accuracy on the CIFAR-10 test dataset and the ImageNet-1K validation dataset.
Flops computation for the Swin Transformers and our revisions uses the function “flops” in the official implementation of [17] [^2] https://github.com/microsoft/Swin-Transformer/blob/main/models/swin_transformer.py. To follow their work, the flops from the Softmax function are omitted. Flops computation for the ViT-B-32 and our revisions uses Python library “ptflops” [^3] https://github.com/sovrasov/flops-counter.pytorch, with manually adding the flops from DCT.
IV-A2 COCO Detection and Segmentation Tasks Implementation Details
Training models for the COCO tasks follows the settings in [17]. Models start from ImageNet-1K pre-trained weights. We use AdamW optimizer [57] for 36 epochs with an initial learning rate of 0.0001, weight decay of 0.05, batch size of 16, 3$\times$ schedule, and multi-scale training [63, 64].
IV-B CIFAR-10 Classification
Training vision transformers from scratch on the CIFAR-10 dataset is notoriously difficult. We demonstrate the benefits of using our DCT initialization and DCT compression for the Swin Transformer Tiny (Swin-T) on from-scratch CIFAR-10 training to show the effectiveness of our methods on the small-scale dataset. Table II presents the performance comparison on the CIFAR-10 test dataset, where the number of parameters only counts trainable parameters.
TABLE II: CIFAR-10 test dataset evaluation. In column “DCT”, Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Models with higher or lower accuracy are highlighted in pink or cyan, respectively.
@@ IV-B CIFAR-10 Classification
Next, we evaluate the DCT-compressed attention with different truncating rates in Swin-T on the CIFAR-10 dataset. Table II presents when we only keep 25% of DCT coefficients, the number of trainable parameters is reduced from 27.53M to 21.45M (22.1% off), while the accuracy only drops 0.05%. If we keep more DCT coefficients, the accuracy even improves. If we keep 75% of DCT coefficients, the accuracy increases from 90.51% to 90.90%. This is because although the number of trainable parameters is reduced, transforming the input data into the frequency domain using DCT can boost the extraction of important features. As a result, fewer trainable parameters are required.
IV-C ImageNet-1K Classification
ImageNet-1K [54] is a large-scale visual database designed for visual object recognition research. It was created to facilitate the development, training, and testing of computer vision models, especially in deep learning. ImageNet consists of a vast collection of labeled images, covering a wide range of object categories.
Table III presents the performance comparison on the ImageNet-1K dataset, where the number of parameters only counts trainable parameters. First, we investigate the DCT-initialized attention in the Swin-T and the Swin-S on the ImageNet-1K dataset. The accuracy on the ImageNet-1K validation dataset is always improved nevertheless which weight matrix is initialized as the DCT matrix. For example, when we initialize the weights to compute keys ($\mathbf{K}$) as the DCT weights, the accuracy of Swin-T from 81.474% to 81.728% and Swin-S are improved from 83.196% to 83.378%, respectively.
Then, we evaluate the DCT-compressed attention with different truncating rates and compare the results with DCFormer [36], as it is also a DCT-based method to reduce the computational cost. Table III presents that when we keep 75% of DCT coefficients, we can obtain a comparable accuracy result on Swin-T and a slightly better accuracy result on Swin-S. Contrarily, the trainable parameters of the Swin-T and the Swin-S are reduced from 28.29M to 25.45M (13.1% off) and from 49.61M to 44.45M (10.4% off), respectively. Flops are reduced from 4.49G to 4.18G (6.9% off) and 8.13G to 8.74G (7.0% off), respectively. If fewer DCT coefficients are kept, the swim transformers can get further savings on parameters and computational overhead. Meanwhile, the accuracy will drop correspondingly. As a comparison, although DCFormer-SW-T/S [36] with $\tau=2$ reduce the FLOPs significantly ($\tau=2$ means 4×4$\times$4 × frequency compression), the performance drops seriously because the DCFormer takes both low-pass and high-pass DCT components equivalent important.
TABLE III: ImageNet-1K validation dataset evaluation. After DCT, Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Our models with higher or lower top-1 accuracy are highlighted in pink or cyan, respectively.
@@ IV-C ImageNet-1K Classification
IV-D Ablation Study
IV-D1 Multiple DCT Initialization
Table IV presents the ablation study about initializing multiple matrices of $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ as the DCT matrix. Although such an initialization still improves the accuracy for Swin-T and Swin-S, the improvement is smaller than with the single DCT initialization. For example, in the Swin-T experiments, if $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$ are both initialized as the DCT matrix, the top-1 accuracy is 81.540%, lower than 81.568% from initializing $\mathbf{W}_{Q}$ only and 81.728% from initializing $\mathbf{W}_{K}$ only. This deterioration is caused by the weight symmetry problem [6] from initializing the multiple weights as the same values. In other words, the attention mechanism is designed to attend to relevant information in the input sequence. If $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the same, the model may not effectively capture the relationships between different positions in the input sequence. Moreover, one of the advantages of self-attention is its ability to model complex dependencies between different positions in the sequence. If $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are the same, the capacity of the model to capture diverse patterns will be limited, and the performance of the model will drop correspondingly.
TABLE IV: ImageNet-1K ablation study on multiple weights initialization as the DCT matrix.
@@ IV-D1 Multiple DCT Initialization
IV-D2 Non-Trainable DCT Weights Initialization
In this case, computing queries, keys, and values is equivalent to shifting the DCT representation of the input patterns. Table V shows we obtain comparable accuracy results as the baseline but use fewer trainable parameters on the CIFAR-10 dataset. It also shows on the ImageNet-1K dataset if we take non-trainable DCT weights to compute $\mathbf{Q}$ or $\mathbf{K}$, the top-1 accuracy drops insignificantly (less than 0.3%). It is worthwhile noting that in this work, the DCT is implemented by matrix multiplication. Hence, the computational overhead of using non-trainable DCT weights is the same as the vanilla attention. Still, the DCT with butterfly operation can further reduce this cost as we only need to apply DCT in each attention layer.
TABLE V: Ablation study on non-trainable DCT weights.
@@ IV-D2 Non-Trainable DCT Weights Initialization
IV-D3 Removing DCT in DCT-Compressed Attention
This section investigates the necessity of DCT in the DCT-compressed attention. As Table VI presents, when we remove the DCT in the DCT-compressed attention and retain other components, the accuracy of the revised Swin-T models on ImageNet-1K decreases across all compression ratios. These experiments show that DCT is essential because it transforms feature representations into the frequency domain, reducing redundancy and enhancing the model’s ability to capture essential information. As a result, the attention mechanism benefits from improved expressiveness and robustness, even under compression. The accuracy drop observed in Table VI confirms that DCT plays a crucial role in preserving discriminative features, demonstrating its necessity in the proposed approach.
TABLE VI: ImageNet-1K ablation study on removing DCT in DCT-compressed attention.
Model Params (M) Top-1 Acc. Top-5 Acc.
Swin-T 28.29 81.474% 95.776%
Swin-T-DCT-0.25 22.21 79.580% 94.936%
Swin-T-DCT-0.25 (Remove DCT) 22.21 78.710% 94.638%
Swin-T-DCT-0.5 23.43 80.762% 95.520%
Swin-T-DCT-0.5 (Remove DCT) 23.43 80.276% 95.086%
Swin-T-DCT-0.75 25.45 81.474% 95.728%
Swin-T-DCT-0.75 (Remove DCT) 25.45 81.064% 95.468%
IV-E COCO 2017 Object Detection and Instance Segmentation
We further evaluate the proposed methods on COCO 2017 object detection and instance segmentation tasks [55]. We use the Swin Transformers with ImageNet-1K pre-trained weights as the backbone and the Cascade Mask R-CNN [65, 66] as the pipeline. Since initializing $\mathbf{W}_{K}$ as the DCT matrix obtains the best accuracy results in the ImageNet-1K experiments, we use Swin-T-DCT-K and Swin-S-DCT-K to evaluate the DCT-based initialization. As Table VII presents, both Swin-T-DCT-K and Swin-S-DCT-K obtain slightly better average precision (AP) results than the vanilla models in the two tasks. Swin-T-DCT-0.75 obtains comparable AP results as Swin-T with about 3M fewer parameters, and Swin-S-DCT-0.75 gets better AP results than Swin-S with about 5M fewer parameters.
TABLE VII: COCO results. Our methods with higher or lower $\text{AP}^{\text{Mask}}$ are highlighted in pink or cyan.
@@ IV-E COCO 2017 Object Detection and Instance Segmentation
V Conclusion
In this work, we introduced two new methods using Discrete Cosine Transform (DCT) to decorrelate the input in the attention for vision transformers. First, we presented a novel DCT-based initialization approach to address the challenges associated with training attention weights. We demonstrated a significant improvement over traditional initialization, providing a robust foundation for the attention mechanism. Second, we formulated a DCT-based compression method that can effectively reduce the computational overhead by truncating high-frequency components. It resulted in smaller weight matrices that compromised considerable accuracy. We validated the efficacy of these breakthroughs, showcasing their potential to enhance the performance and efficiency of Transformer models in the classification tasks. These contributions underscore the importance of thoughtful initialization strategies and compression techniques in advancing the capabilities of the Swin Transformer architectures.
VI Acknowledgment
This work was supported by National Science Foundation (NSF) grants 1934915, 2229659, U.S. Department of Energy (DOE) grant DE-SC0023715, and National Institutes of Health (NIH) grant R01-CA240639.